OpenStudy (anonymous):

I'm having some issues with integration by u-substitution. Solving in general, but also how can I determine a u-value that isn't just guess and check? Here's the problem. Integrate(3(dx))/(x(ln(7x)))

OpenStudy (jack1):

because these are all " ax + bx + c", u dont really need to worry about choosing a u substitution... if it becomes ax / (2+bx) +c ... then start worrying about u values When i do it, if there's times or divide involved in the function im integrating, then i start to worry about substitution, if it's simply addition and or subtraction... i dont bother

OpenStudy (jack1):

so this function id'd do sequentially: integral of 6x^5 = ??

OpenStudy (anonymous):

I know how to integrate this without u-substitution, but this one specifically asked for it to be solved in that manner.

OpenStudy (anonymous):

To answer your question, \[\int\limits 6x ^{5}=x ^{6}\]

OpenStudy (jack1):

weird... there's no need for chain, product or quotient rule here...

OpenStudy (jack1):

i guess you could substitute u = 6x into the eqn... if you really have to... ???

OpenStudy (anonymous):

I'll get a different example. \[\int\limits \frac{ 3dx }{ x(\ln(7x)) }\]

hartnn (hartnn):

for your first example, after you would have separated out the integral, your first term will be \(\int 6x^5dx\) now if you "have to" use substitution, you would put \(\large u= x^6\) du =6x^5 dx \(\implies \int u du\) and after integrating , substitute back u = x^6 for 2nd, you can just use u=ln(7x)

hartnn (hartnn):

***sorry, you would get \(\large \int du\)

hartnn (hartnn):

jack! whats your u ?? and du ?

OpenStudy (anonymous):

I believe I should choose u=ln(7x) which would make\[du=\frac{ 1 }{ 7x }(7)=\frac{ 1 }{ x }\]

hartnn (hartnn):

whats your du then ?? you missed a point, take u as ln (7x) only, just as brenar did

hartnn (hartnn):

yes, that would lead to \(\large \int 3du\) easy to integrate, right! thats why we use substitution method in integration....

hartnn (hartnn):

**\(\large \int (3/u) du\)

OpenStudy (anonymous):

I guess I get lost after determining a u and du. I can substitute in, giving me \[\int\limits \frac{ 3dx }{ xu }\] I know that the above statement is incorrect because I've substituted in u, which means I can't have dx in there, but... I'm not sure where to go from here.

OpenStudy (jhannybean):

Integration by parts?

hartnn (hartnn):

you already know, that du = (1/x) dx , right ?? \(\huge \int\limits \frac{ 3dx }{ xu } = \int\limits \dfrac{ 3 }{ u }\times \dfrac{dx}{x}=\int\limits \dfrac{ 3 }{ u }\times du\) got this ?

OpenStudy (jhannybean):

@hartnn how do you separate out this integral?...

OpenStudy (anonymous):

Apologies for being rather thick headed, but where did you get \[\int\limits \frac{ 3 }{ u } \times du\] I understand how you separated my original integral, but not the simplification it would seem.

hartnn (hartnn):

u= ln 7x du/dx = 1/ (7x) d/dx (7x) = 1/x so, du = (1/x) dx got this ?

OpenStudy (anonymous):

Yes.

hartnn (hartnn):

jhanny, whats your doubt exactly ?

hartnn (hartnn):

so, 3dx/ xu = (3/u) (1/x) dx = (3/u) du this ?

OpenStudy (jhannybean):

im with you on this, caught up, lol.

OpenStudy (anonymous):

(3/u) (1/x) dx = (3/u) du Where did the dx go?

hartnn (hartnn):

nowhere "u= ln 7x du/dx = 1/ (7x) d/dx (7x) = 1/x so, du = (1/x) dx" (3/u) (1/x) dx = (3/u) du

OpenStudy (jhannybean):

omg that's way too cool.

hartnn (hartnn):

brenar, i think you have in mind that, du = 1/x its actually, du = (1/x) dx

OpenStudy (anonymous):

\[\int\limits \frac{ 3 }{ xln(7x) }=3\ln \left| \ln(7x) \right|\]

hartnn (hartnn):